OpenStudy (anonymous):

4b/2 + 3 = 11 + b It's amazing what one would forget in a year without a math class. If someone could show it well step-by-step, that would be ideal. Thanks!

OpenStudy (amistre64):

clear out the fraction by multiplying it all by 2

OpenStudy (amistre64):

i assume you are trying to solve for b?

OpenStudy (amistre64):

another option is to see that 4/2 = 2 and go from that :)

OpenStudy (amistre64):

when working with equations; always remember to do the same thing to each side in order to keep them balanced

OpenStudy (anonymous):

I don't know why multiplying it all by two didn't cross my mind. I got: \[4b + 6 = 22+ 2b\] \[2b + 6 = 22\] \[2b = 16\] \[b = 8\] Thank you much!

OpenStudy (amistre64):

that looks to be good :)